OpenStudy (anonymous):

1/3(6x - 12) = 4(1/2x + 1) -2

OpenStudy (precal):

\[(1/3)(6x-12)=4((1/2)x+1)-2\] 2x-4=2x+4-2 2x-4=2x+2 -4=2 No Solution